OpenStudy (anonymous):

simplify (x^2-1/9x)/(x^2+2x+1/3x^2)

OpenStudy (campbell_st):

factorise the fractions 1st \[\frac{(x+1)(x-1)}{3 \times 3x}\times \frac{3x\times x}{(x+1)^2}\] cancel common factors leaves \[\frac{x(x-1)}{3(x+1)}\]

OpenStudy (anonymous):